### v2

@@ 4.3 Specific application: laser cutting
TRUMPF laser cutting machines possess the capability to operate autonomously once the desired cutting program is specified, handling everything from running the laser path to loading or unloading and sorting materials. This enables the machine to operate without human supervision, yet frequent production standstills persist due to thermal expansion on laser-cut parts, causing up to 1% of parts getting stuck inside the sheet and having to be removed manually. Consequently, laser cutting machines can run into a standstill overnight, which results in significant economic disadvantage. To solve this problem, it is proposed to simulate the heat distribution during the laser cutting process. By accurately modeling it and subsequently predicting thermal expansion, an optimized cutting route that minimizes the risk of machine standstills can be planned.
Accurately simulating the temperature during laser cutting with complex geometries and long cutting times requires significant computational efforts. In contrast to conventional FEM-based simulations, ML approaches, such as MeshGraphNets pfaff2020mgn , accelerate FEM simulations by one to two orders of magnitude whilst maintaining a high accuracy. However, training of such large ML models currently takes several weeks, which makes updating such models computationally challenging. In this work, the advantages of quantum machine learning approaches, as described above, are investigated to simulate thermodynamics from laser-induced heat on metal plates.
The metal geometry is represented by a mesh structure which can be formulated as a graph $G=(V,E)$. The model is formulated employing the PQC approach outlined in section 4.2 for predicting the temperature $f_{v}^{t+1}$ at the vertex $v$ for the next time step $(t+1)$, depending on the temperatures from the previous time step $t$ at the vertex $v$ and its $d$ neighbors $\mathcal{N}(v)$. A visualisation of the circuit can be seen in Fig. 3.

### v3

@@ 4.3 Specific application: laser cutting
TRUMPF laser cutting machines possess the capability to operate autonomously once the desired cutting program is specified, handling everything from running the laser path to loading or unloading and sorting materials. This enables the machine to operate without human supervision, yet frequent production standstills persist due to thermal expansion on laser-cut parts, causing up to 1% of parts getting stuck inside the sheet and having to be removed manually. Consequently, laser cutting machines can run into a standstill overnight, which results in significant economic disadvantage. To solve this problem, it is proposed to simulate the heat distribution during the laser-cutting process. By accurately modeling it and subsequently predicting thermal expansion, an optimized cutting route that minimizes the risk of machine standstills can be planned.
Accurately simulating the temperature during laser cutting with complex geometries and long cutting times requires significant computational efforts. In contrast to conventional FEM-based simulations, ML approaches, such as MeshGraphNets pfaff2020mgn , accelerate FEM simulations by one to two orders of magnitude whilst maintaining a high accuracy. However, training such large ML models currently takes several weeks, which makes updating such models computationally challenging. Efforts have been made to conduct thermal simulations using hybrid-QML approaches, such as QFE-GNNs XuSascha2024QuantumFE , where the encoding layer of the classical GNN is substituted with PQC to enhance feature representation. This method achieves better performance than traditional GNNs, however, scaling these models to handle larger graphs with complex geometries remains challenging. In this work, the advantages of quantum machine learning approaches, as described above, are investigated to simulate thermodynamics from laser-induced heat on metal plates.
The metal geometry is represented by a mesh structure, which can be formulated as a graph $G=(V,E)$. The model is formulated employing the PQC approach outlined in section 4.2 for predicting the temperature $f_{v}^{t+1}$ at the vertex $v$ for the next time step $(t+1)$, depending on the temperatures from the previous time step $t$ at the vertex $v$ and its $d$ neighbors $\mathcal{N}(v)$. A visualization of the circuit can be seen in Fig. 3.